The engine made by alpha-x for boats have an average power of 220 horsepower and a standard deviation of 15HP. You can assume the distribution of power follows a normal distribution. Beta-y is testing the engines and will dispute the company's claim if the sample mean is less than 215HP. If they take a sample of 4 engines; what is the probability the mean is less than 215? If beta-y samples 100 engines, what is the probability that the mean will be less than 215?


1
Expert's answer
2021-11-12T11:13:34-0500

Solution:

We want to find "P(\bar X<215)".

Since the population follows a normal distribution, we can conclude that "\bar X"  has a normal distribution with mean 220 HP (μ=220) and a standard deviation of "\dfrac σ{\sqrt n}=\dfrac{15}{\sqrt4}=7.5" HP.

Now,

"P(\bar X<215)=P(z<\dfrac{215-220}{7.5})
\\=P(z<-0.67)
\\\approx 0.2514"

Next, the sampling distribution of the sample mean is Normal with mean μ=220

 and standard deviation "\dfrac σ{\sqrt n}=\dfrac{15}{\sqrt{100}}=1.5"

Now,

"P(\bar X<215)=P(z<\dfrac{215-220}{1.5})
\\=P(z<-3.333..)
\\\approx 0.00043"